MONT 104N – Modeling the Environment Student-Led Discussions November 30 – December 5

General Information

For (roughly) the final week of the semester, different assigned groups from the class will be responsible for giving oral presentations, including some questions to get a class discussion started. The choice of topic will be up to you, but I think you’ll probably want to choose one of the articles you reported in your journals earlier in the semester, or something similar you might have read more recently. Suggestions:

• There are a number of interesting aspects of the recent devastating California wildfires that might form good topics.

• There have also been several recent stories about effects of climate change on various natural habitats.

• Scientists have also very recently announced the discovery of a massive asteroid impact crater located under the Greenland ice sheet and used evidence they found there to deduce features of the climate in Greenland before the ice sheet formed.

Important Note: Please notify me (either by email or by coming to office hours to discuss your ideas) about the topic you want to present no later than Monday, November 26 (the day we return from Thanksgiving break). I reserve the right to ask you to think about a different topic or to come at a topic from a different angle if two groups end up asking to do the same thing.

Let me know if you need any help tracking down sources for these or other possible topics.

The assignment will be:

• To select a timely news story involving some environmental question,

• To present what the story says, possibly including connections with things we have talked about in the course (if they exist), and then

• To get a class discussion started about the issues involved, what people are trying to do about it, what our responses should be, and so forth.

Each group will have about half a class period (25 minutes). Aim to use about half of your group’s time presenting what your source story or article said, and the rest on the followup discussion. You may use PowerPoint or GoogleSlides (which is similar to PowerPoint) for the presentation segment if you like. If you choose that option, you probably don’t want any more than 5 slides, though, so choose carefully what you put there. As a general guideline, I would say the slides should be used to present visual information (i.e. images, graphs, charts, ... ) rather than to duplicate what you are saying. Your oral presentation will explain what’s going on, tell us what to look for, and so forth. Please don’t copy and paste whole pages from your news story or article into the slides you present.

These presentations will be going on in the class meetings of Friday, November 30, Monday, December 3 (after the first 15 minutes or so, when the CEFs – course evaluation forms – for MONT 104N will be administered), Wednesday, December 5, and possibly on Friday, December 7, depending on how long things go in the previous days. That Friday, our last class meeting this semester, I will also let you know about the format and what to expect about the content of the final exam. Schedule Please take a look at this before you leave for Thanksgiving and let me know if there are any conflicts or situations that might call for changes – I’m willing to consider rearranging things as long as everyone involved is OK with it. November 30 Group 1 – Kerry, Ryan, Courtney Group 2 – Tess, Carlie, Joe December 3 Group 3 – Shane, Aidan, Shanis December 5 Group 4 – Rebecca, Ben, Nick Group 5 – Clarisa, Ezra, Will, Jack